# HSG-12M: A Large-Scale Benchmark of Spatial Multigraphs from the Energy Spectra of Non-Hermitian Crystals

### v2

@@ 1 Introduction
A rich resource lies in the Hamiltonian spectral graph, a fascinating and diverse object emerging from recent advances in non-Hermitian physics. Recent advances have shown that the energy spectrum of one-dimensional crystals under open boundary conditions [^3] To be precise, it is 1-D crystal (lattice) Hamiltonian, under open boundary conditions (OBC), in the thermodynamic limit (i.e. large-size limit, the length of the lattice $\to\infty$). forms arcs and loops on the complex energy plane. These spectral loci can be naturally represented as spatial graphs embedded in the two-dimensional $\mathcal{C}$}-plane. Moreover, these spectral graphs (Tai & Lee, 2023; Lin et al., 2023; Xiong & Hu, 2023; Wang et al., 2024) serve as fingerprints with far more intricate structures than conventional topological signatures for electronic behavior (e.g., $\mathbb{Z}/\mathbb{Z}_{2}$ invariants, Chern number (Hasan & Kane, 2010)). Figure 2&A4 show examples of these graphs, featuring a kaleidoscope of nontrivial edge geometries and multiplicities that form diverse patterns beyond existing graph datasets.
@@
Figure 1: Number of graphs v.s. number of classes in HSG-12M compared to other graph-classification datasets. HSG-12M is the only large-scale multigraph (i.e. unlike simple graph that only allows one edge between any node pair) dataset, with exceptional class diversity even exceeds all other simple graph datasets. T-HSG-5M holds temporal spatial multigraphs. Table A4 lists comprehensive comparison against 45 other datasets.
@@ 2 Poly2Graph: Automating Spectral Graph Extraction
Open-Source Release and Broader Impact. Poly2Graph1 is released under the MIT licence. See a tutorial in appendix G. Poly2Graph establishes a turn-key mechanism for translating linear operators into machine-learning-ready graphs, bridging condensed matter physics and graph representation learning. The same principle extends to any vector, matrix, and univariate/bivariate polynomial, opening an new “algebra-as-graph” perspective, broadening the applicability of Poly2Graph to a wide range of other areas (section 5, appendix F).
@@ 3 HSG-12M Dataset Description
Data Format and Accessibility. To maximize accessibility and flexibility, we release HSG-12M under a permissive CC BY 4.0 license. The dataset is publicly available via Dataverse (Yan et al., 2025). Users can download the full dataset or select specific subsets using the code provided at https://github.com/sarinstein-yan/HSG-12M. In companion, we release an auxiliary package HSG-12M2 for effortless data handling, benchmark reproduction, custom featurization and dataset generation, interactive tutorial, and more.
@@ D.1 Comparison with 45 Other Datasets
Table A4: Overview of graph-level benchmark datasets. Each row corresponds to one dataset: #Graphs gives the total number of graphs; #Classes is the number of target labels; #Nodes and #Edges report the average number of nodes and edges per graph; Scale column follows the OGB Large-Scale Challenge definitions [37]: Small (S) datasets fall below the large-scale thresholds; Medium (M) datasets contain >> 1 million nodes or >> 10 million edges; Large (L) datasets exceed 100 million nodes or 1 billion edges.; Attributed indicates whether both node and edge features are present; Spatial denotes whether the graphs carry geometric or coordinate information (e.g. 2D, 3D, geographic coordinate system-GCS); Temporal flags static (S) or time‐series graph data (T); Multi marks support for multiple edges between node pairs; and RCG (Rich Connection Geometry) indicates datasets whose edge geometry exhibits nontrivial patterns that go beyond simple straight-line connections. “?” entries indicate information not stated in the original paper. In addition to values extracted from the literature, some benchmark statistics were sourced from Refs. [37; 98; 21; 66].
@@
MUTAG [50; 63] 188 2 17.9 19.8 S S
SIDER [88; 1] 1.4K 2 33.6 35.4 S S
BACE [88; 76] 1.5K 2 34.1 36.9 S S
ClinTox [88; 28] 1.5K 2 26.2 27.9 S S
BBBP [88; 60] 2.0K 2 24.1 25.9 S S
Tox21 [88; 80] 7.8K 2 18.6 19.3 M S
ToxCast [88; 70] 8.6K 2 18.8 19.3 M S
@@
IMDB-BINARY [63; 95] 1 K 2 19.8 96.5 S S
IMDB-MULTI [63; 95] 1.5 K 3 13.0 65.9 S S
REDDIT-BINARY [63; 95] 2 K 2 429.6 497.8 S S
REDDIT-MULTI-5K [63; 95] 5.0 K 5 508.5 594.9 M S
REDDIT-MULTI-12K [63; 95] 11.9 K 11 11.0 391.4 M S
@@ E.5 Physical Motivation of the Benchmark Task
Beyond enabling inverse design, the HSG provides a large-scale resource to study how Hamiltonian parameters control spectral-graph morphology (Yan et al., ) and to pre-train scientific foundation models on physically grounded graph signals. In short, the mapping $\texttt{Spectral Graph}\!\to\!\texttt{ChP Class}$ directly operationalizes a task that materials physicists and chemists already perform by hand, but at scale and with principled uncertainty via Top-$k$ retrieval.

### v3

@@ 1 Introduction
A rich resource lies in the Hamiltonian spectral graph, a fascinating and diverse object emerging from recent advances in non-Hermitian physics. Recent advances have shown that the energy spectrum of one-dimensional crystals under open boundary conditions [^3] To be precise, it is 1-D crystal (lattice) Hamiltonian, under open boundary conditions (OBC), in the thermodynamic limit (i.e. large-size limit, the length of the lattice $\to\infty$). forms arcs and loops on the complex energy plane. These spectral loci can be naturally represented as spatial graphs embedded in the two-dimensional complex plane ($\mathbb{C}$}-plane). Moreover, these spectral graphs (Tai & Lee, 2023; Lin et al., 2023; Xiong & Hu, 2023; Wang et al., 2024) serve as fingerprints with far more intricate structures than conventional topological signatures for electronic behavior (e.g., $\mathbb{Z}/\mathbb{Z}_{2}$ invariants, Chern number (Hasan & Kane, 2010)). Figure 2&A4 show examples of these graphs, featuring a kaleidoscope of nontrivial edge geometries and multiplicities that form diverse patterns beyond existing graph datasets.
@@
Figure 1: Number of graphs v.s. number of classes in HSG-12M compared to other graph-classification datasets. HSG-12M is the only large-scale multigraph (i.e. unlike simple graph that only allows one edge between any node pair) dataset, with exceptional class diversity that even exceeds all other simple graph datasets. T-HSG-5M holds temporal spatial multigraphs. Table A4 lists comprehensive comparison against 45 other datasets.
@@ 2 Poly2Graph: Automating Spectral Graph Extraction
Open-Source Release and Broader Impact. Poly2GraphLABEL:pkg:p2g is released under the MIT licence. See a tutorial in appendix G. Poly2Graph establishes a turn-key mechanism for translating linear operators into machine-learning-ready graphs, bridging condensed matter physics and graph representation learning. The same principle extends to any vector, matrix, and univariate/bivariate polynomial, opening an new “algebra-as-graph” perspective, broadening the applicability of Poly2Graph to a wide range of other areas (section 5, appendix F).
@@ 3 HSG-12M Dataset Description
Data Format and Accessibility. To maximize accessibility and flexibility, we release HSG-12M under a permissive CC BY 4.0 license. The dataset is publicly available via Dataverse (Yan et al., 2025a). Users can download the full dataset or select specific subsets using the code provided at https://github.com/sarinstein-yan/HSG-12M. In companion, we release an auxiliary package HSG-12MLABEL:pkg:hsg for effortless data handling, benchmark reproduction, custom featurization and dataset generation, interactive tutorial, and more.
@@ Appendix A Related Work
A.4 Non-Bloch Band Theory
Non-Hermitian band theory has developed a mature language for classifying complex spectra and their topological phases, including point-/line-gap notions, symmetry-based classifications, and associated boundary phenomena (Lee & Thomale, 2019; Kawabata et al., 2019; Gong et al., 2018; Li & Lee, 2022). A central theme is the breakdown (and restoration) of bulk–boundary correspondence via non-Bloch descriptions, which underpin the non-Hermitian skin effect (Kunst et al., 2018; Jin & Song, 2019; Yang et al., 2020; Lin et al., 2023; Xiong et al., 2024). Recent advances further sharpen the geometric viewpoint on complex band structures through generalized Brillouin-zone constructions and higher-dimensional non-Bloch theory, e.g., the amoeba formulation (Wang et al., 2024; 2025b; Kim et al., 2025; Meng et al., 2026). In parallel, there is growing interests on “spectral geometry” and morphology in the complex plane—from graph-topological characterizations of spectra and band morphology to the structure of self-intersections and spectral transitions (Tai & Lee, 2023; Xiong & Hu, 2023; Yan et al., 2025b; Pi et al., 2025; Zhong et al., 2025). Real-space geometry dependence also appears in experimental settings and in general theories (Fang et al., 2022; Wang et al., 2025a). Beyond classification, there exists dynamical and response-based viewpoints (Li et al., 2021; Lee et al., 2020; Xu et al., 2025). Finally, a growing experimental and digital-quantum literature has begun to directly probe these effects (Wang & Chen, 2025; Zhang et al., 2025; Shen et al., 2025; Koh et al., 2025; Yang et al., 2022).
@@ D.1 Comparison with 45 Other Datasets
Table A4: Overview of graph-level benchmark datasets. Each row corresponds to one dataset: #Graphs gives the total number of graphs; #Classes is the number of target labels; #Nodes and #Edges report the average number of nodes and edges per graph; Scale column follows the OGB Large-Scale Challenge definitions [39]: Small (S) datasets fall below the large-scale thresholds; Medium (M) datasets contain >> 1 million nodes or >> 10 million edges; Large (L) datasets exceed 100 million nodes or 1 billion edges.; Attributed indicates whether both node and edge features are present; Spatial denotes whether the graphs carry geometric or coordinate information (e.g. 2D, 3D, geographic coordinate system-GCS); Temporal flags static (S) or time‐series graph data (T); Multi marks support for multiple edges between node pairs; and RCG (Rich Connection Geometry) indicates datasets whose edge geometry exhibits nontrivial patterns that go beyond simple straight-line connections. “?” entries indicate information not stated in the original paper. In addition to values extracted from the literature, some benchmark statistics were sourced from Refs. [39; 119; 21; 78].
@@
MUTAG [56; 75] 188 2 17.9 19.8 S S
SIDER [105; 1] 1.4K 2 33.6 35.4 S S
BACE [105; 90] 1.5K 2 34.1 36.9 S S
ClinTox [105; 29] 1.5K 2 26.2 27.9 S S
BBBP [105; 71] 2.0K 2 24.1 25.9 S S
Tox21 [105; 94] 7.8K 2 18.6 19.3 M S
ToxCast [105; 83] 8.6K 2 18.8 19.3 M S
@@
IMDB-BINARY [75; 116] 1 K 2 19.8 96.5 S S
IMDB-MULTI [75; 116] 1.5 K 3 13.0 65.9 S S
REDDIT-BINARY [75; 116] 2 K 2 429.6 497.8 S S
REDDIT-MULTI-5K [75; 116] 5.0 K 5 508.5 594.9 M S
REDDIT-MULTI-12K [75; 116] 11.9 K 11 11.0 391.4 M S
@@ E.5 Physical Motivation of the Benchmark Task
Beyond enabling inverse design, the HSG provides a large-scale resource to study how Hamiltonian parameters control spectral-graph morphology and to pre-train scientific foundation models on physically grounded graph signals, contributing to AI4Science community (Yan, 2025; Yan & Pan, 2025). In short, the mapping $\texttt{Spectral Graph}\!\to\!\texttt{ChP Class}$ directly operationalizes a task that materials physicists and chemists already perform by hand, but at scale and with principled uncertainty via Top-$k$ retrieval.